OpenStudy (anonymous):

Divide.Simplify.4/5 divided by 7/10

OpenStudy (anonymous):

Find the lowest common denominator and divide.

OpenStudy (anonymous):

how do i find the lowest common denominator?

OpenStudy (anonymous):

What is the lowest denominator that both can be? So for this, it's 10 so multiply \[\frac{ 4 }{ 5 }\] by 2

OpenStudy (anonymous):

4/5 x 2 = 8/10 8/10 divided by 7/10 Both denominators are crossed out so 8/7 = 1 and 1/7

OpenStudy (anonymous):

ARIGATO=THANK YOU IN JAPAN